OpenStudy (anonymous):

Is the equation an example of growth or decay? y = 2.1(0.15)x A. growth B. decay

OpenStudy (anonymous):

B decay since the common ratio is less than 1

OpenStudy (paxpolaris):

y = 2.1(0.15)^x \[y = 2.1(0.15)^x\]